Maximum value of cos^3x+cos^3(120+x)+cos^3(120-x)

Grade:11

1 Answers

Sher Mohammad IIT Delhi
9 years ago
cos^3x+cos^3(120+x)+cos^3(120-x)=f(x)
f'(x)= 3cos2xsinx+3cos2 (x+120)sin(x+120)+
3cos2(x-120)sin(x-120) =0
solve for x to get the maxima
sher mohammad
b.tech, iit delhi